The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
Answer by checkley71(8403)

THE PERIMETER IS =B+L1+L2 L1 & L2 = 4B-8 THUS
B+4B-8+48-8=29
9B=29+16
9B=45
B=45/9
B=5 FOR THE BASE THUS THE SIDE IS
4*5-8=20-8=12
PROOF
5+12+12=29
29=29